OpenStudy (anonymous):

A student studying for a vocabulary test knows the meanings of 12 words from a list of 20 words. If the test contains 9 words from the study list, what is the probability that at least 7 of the words on the test are words that the student knows?

OpenStudy (anonymous):

@mathmale can you help me please?

OpenStudy (anonymous):

@ganeshie8 can you help me please?

ganeshie8 (ganeshie8):

looks like a binomial distribution, but not entirely sure... @amistre64

ganeshie8 (ganeshie8):

P(know) = 12/20 = 3/5 P(dont know) = 8/20 = 2/5

ganeshie8 (ganeshie8):

the test has 9 words, P(X >= 7) = P(X = 7) + P(X = 8) + P(X = 9)

ganeshie8 (ganeshie8):

does that look familiar ha ?

OpenStudy (anonymous):

No :\